OpenStudy (anonymous):

A person of mass 60 kg is inside a lift of mass 940 kg and presses the button on control panel. The lift starts moving upwards with an acceleration 1.0 m/s2. If g510 ms22, calculate the tension in the supporting cable.

OpenStudy (anonymous):

Um, what exactly does "g510 ms22" mean?

OpenStudy (anonymous):

 I'm guessing that meant the question wants you to use a g value of 10m/s^2. Anway, F = ma. In this case, F = T. T - w = ma. T = ma + mg. T = (1000kg)(1m/s^2) + (1000kg)(10m/s^2). T = 1.1x10^4 N.